# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 3.1. Riemannian metrics from the family
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},(\mathfrak{a})_{\mathrm{skew}}=\{0\}$. In this case, $\mathfrak{g}\subset\mathcal{E}_{\mathrm{sym}}$, an example is a group of commutative symmetric matrices.
A generalization of theorem 1 is to replace $\beta_{1}$ with a symmetric operator $\mathrm{B}_{1}$ on $\mathfrak{a}$, such that $[\mathrm{B}_{1}\mathtt{a},\mathtt{b}]=\mathrm{B}_{1}[\mathtt{a},\mathtt{b}]$ for $\mathtt{a},\mathtt{b}\in\mathfrak{a}$ ($\mathrm{B}_{1}$ is called an intertwining operator), such that $\mathrm{B}_{1}$ has nonzero eigenvalue, we can check all the arguments in the proof of theorem 1 still hold. For $\mathfrak{so}(n)$ (a simple Lie algbra), such an operator must be a scalar multiplication. If $\mathfrak{a}$ is composed of several simple components, $\mathrm{B}_{1}$ could be constructed by applying different multipliers on different components. We discuss this briefly in section 4.
4. The Generalized Linear group with positive determinant
In this case, the group is $\mathrm{G}=\operatorname{GL}^{+}(n)$, the group of invertible matrices with positive determinant in $\mathbb{R}^{n\times n}$, with Lie algebra $\mathfrak{g}=\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$. We take $\mathfrak{a}=\mathfrak{so}(n)$, the Lie algebra of antisymmetric matrices. For $\mathtt{g}\in\mathfrak{g}$, we have $\mathtt{g}_{\mathfrak{a}}=\mathtt{g}_{\mathrm{skew}}$. We typically take $\beta_{0}=1$, for $\mathtt{g}\in\mathfrak{g}$, the metric becomes
(28) $\operatorname{Tr}\mathtt{g}^{2}-(\beta+1)\operatorname{Tr}\mathtt{g}_{\mathrm{skew}}\mathtt{g}_{\mathrm{skew}}=\operatorname{Tr}\mathtt{g}^{2}_{\mathrm{sym}}+\beta\operatorname{Tr}\mathtt{g}_{\mathrm{skew}}^{\mathsf{T}}\mathtt{g}_{\mathrm{skew}}=\lvert\mathtt{g}\rvert_{\mathcal{F}}^{2}+(\beta-1)\lvert\mathtt{g}_{\mathrm{skew}}\rvert_{\mathcal{F}}^{2}$
after the orthogonal decomposition $\mathtt{g}=\mathtt{g}_{\mathrm{sym}}+\mathtt{g}_{\mathrm{skew}}$. When $\beta=1$, this is the Frobenius form, covered in [28]. Expanding the $\mathrm{skew}$ operator, theorem 1 implies
@@ 4. The Generalized Linear group with positive determinant
The geodesic starting at $X$ with initial velocity $\xi$ is given by
(30) $\gamma(t)=X\exp(\frac{t}{2}\left((1-\beta)\xi+(1+\beta)\xi^{\mathsf{T}}\right))\exp(t(1+\beta)\xi_{\mathrm{skew}}).$
@@
(31) $\mathtt{T}_{0}^{t}(\gamma)\eta=X\exp(\frac{t}{2}\left((1-\beta)\xi+(1+\beta)\xi^{\mathsf{T}}\right))\operatorname{expv}(t\mathrm{P},X^{-1}\eta)\exp(t(1+\beta)(\xi_{\mathrm{skew}})),$
where the operator $\mathrm{P}=\mathrm{P}_{\mathtt{a}}$ for $\mathtt{a}=X^{-1}\xi$ is given by
(32) $\mathrm{P}_{\mathtt{a}}\mathtt{b}=\frac{1}{2}([\mathtt{b},\mathtt{a}]+(1+\beta)([\mathtt{a}_{\mathrm{skew}},\mathtt{b}]-[\mathtt{b}_{\mathrm{skew}},\mathtt{a}])).$
@@
Instead of taking $\mathfrak{a}=\mathfrak{so}(n)$, if we take $\mathfrak{a}=\mathfrak{so}(n)\oplus\mathbb{R}\operatorname{I}_{n}$, then this violates condition item 3. However, if instead of taking a scalar $\beta_{1}$, we take the operator $\mathrm{B}_{1}$ which acts by multiplying $\mathfrak{so}(n)$ by $\beta_{1}$ and $\mathbb{R}\operatorname{I}_{n}$ by a negative constant, so that the metric in section 3.1 is Riemannian, the formulas still hold, as $\mathbb{R}\operatorname{I}_{n}$ does not contribute to the Lie brackets. This is the metric in [18].
The case $\beta=-1$ corresponds to the trace form and the pseudo-Riemannian Khvedelidze-Mladenov metric [14].
5. The Special Orthogonal group with a deformed metric
@@
In this case, $\mathrm{G}=\mathrm{SO}(n)$, the group of orthogonal matrices with determinant $1$. Its Lie algbra $\mathfrak{g}=\mathfrak{so}(n)$ is identified with the set of antisymmetric matrices. For $\mathfrak{a}$, we take the algebra of matrices where the top $d\times d$ block is antisymmetric and the remaining entries vanish, ($\mathfrak{a}=\begin{bmatrix}\mathfrak{so}(d)&0_{d\times(n-d)}\\ 0_{(n-d)\times d}&0_{(n-d)\times(n-d)}\end{bmatrix}$ graphically). Here, the projection $\mathrm{p}_{\mathfrak{a}}$ from $\mathbb{R}^{n\times n}$ maps a matrix $\omega$ to the matrix $\omega_{\mathfrak{a}}$ with the top $d\times d$ block is antisymmetrized from the corresponding block of $\omega$, while the remaining entries vanish. The standard normalization takes $\beta_{0}=-\frac{1}{2}$. Setting $\beta_{1}=\alpha$, we have
@@
At $X\in\mathrm{SO}(n)$, $\xi,\eta\in T_{X}\mathrm{SO}(n)$ where $\xi=X\begin{bmatrix}A_{\xi}&-B^{\mathsf{T}}_{\xi}\\ B_{\xi}&C_{\xi}\end{bmatrix}$, $\eta=X\begin{bmatrix}A_{\eta}&-B^{\mathsf{T}}_{\eta}\\ B_{\eta}&C_{\xi}\end{bmatrix}$, with $A_{\xi},A_{\eta}\in\mathfrak{so}(d),C_{\xi},C_{\eta}\in\mathfrak{so}(n-d)$, the metric in theorem 1 with $\beta_{0}=-\frac{1}{2},\beta_{1}=\alpha$ is given by
(33) $\langle\xi,\eta\rangle=\frac{1}{2}\operatorname{Tr}\xi^{\mathsf{T}}\eta+(\alpha-\frac{1}{2})\operatorname{Tr}((\xi^{\mathsf{T}}X)_{\mathfrak{a}}(X^{\mathsf{T}}\eta)_{\mathfrak{a}})=\frac{1}{2}\operatorname{Tr}\xi^{\mathsf{T}}\eta+(\alpha-\frac{1}{2})\operatorname{Tr}A_{\xi}^{\mathsf{T}}A_{\eta}.$
@@
(34) $\begin{gathered}\Gamma(\xi,\eta)=\frac{1}{2}X(\xi^{\mathsf{T}}\eta+\eta^{\mathsf{T}}\xi)+\frac{1-2\alpha}{2}X([(X^{\mathsf{T}}\xi)_{\mathfrak{a}},X^{\mathsf{T}}\eta]+[(X^{\mathsf{T}}\eta)_{\mathfrak{a}},X^{\mathsf{T}}\xi])\\ =\frac{1}{2}X(\xi^{\mathsf{T}}\eta+\eta^{\mathsf{T}}\xi)+\frac{1-2\alpha}{2}X\begin{bmatrix}0_{d\times d}&-A_{\xi}B_{\eta}^{\mathsf{T}}-A_{\eta}B_{\xi}^{\mathsf{T}}\\ -B_{\eta}A_{\xi}-B_{\xi}A_{\eta}&0_{(n-d)\times(n-d)}.\end{bmatrix}.\end{gathered}$
@@
(35) $\gamma(t)=X\exp\left(t\begin{bmatrix}2\alpha A_{\xi}&-B_{\xi}^{\mathsf{T}}\\ B_{\xi}&C_{\xi}\end{bmatrix}\right)\exp\left(t\begin{bmatrix}(1-2\alpha)A_{\xi}&0\\ 0&0\end{bmatrix}\right).$
@@
(36) $\mathtt{T}(\gamma)_{X}^{\mathsf{T}}\eta=X\exp\left(t\begin{bmatrix}2\alpha A_{\xi}&-B^{\mathsf{T}}_{\xi}\\ B_{\xi}&C_{\xi}\end{bmatrix}\right)\operatorname{expv}(t\mathrm{P},X^{\mathsf{T}}\eta)\exp\left(t\begin{bmatrix}(1-2\alpha)A_{\xi}&0\\ 0&0\end{bmatrix}\right)$
where the operator $\mathrm{P}=\mathrm{P}_{\mathtt{a}}$ for $\mathtt{a}=X^{\mathsf{T}}\xi$ is given in eq. 21 with $\beta=-2\alpha$.
6. A quotient metric
The two standard metrics on the Stiefel manifold are instances of the quotient metric defined in this section. This metric takes advantage of the fact that the trace form is bi-invariant, to construct a right-invariant group action of a simple format.
Let $\mathrm{K}\subset\mathrm{G}$ be a connected Lie subgroup. We have an action of $\mathrm{K}$ on $\mathrm{G}$ by right multiplication, the action is $(X,K)\mapsto XK\in\mathrm{G}$, for $X\in\mathrm{G},K\in\mathrm{K}$. We assume the action is proper (a condition that is satisfied if $\mathrm{K}$ is compact, that is the case for the examples we consider here). Then $\mathcal{M}=\mathrm{G}/\mathrm{K}$ has the structure of a quotient manifold. We will show in theorem 2 below that with an additional condition, $\mathrm{K}$ acts on $\mathrm{G}$ as an isometry group of the metric in theorem 1 by right multiplication, and we can apply the theory of pseudo-Riemannian submersion.

### v2

@@ 4.3. Riemannian metrics from the family
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},(\mathfrak{a})_{\mathrm{skew}}=\{0\}$. In this case, $\mathfrak{g}\subset\mathcal{E}_{\mathrm{sym}}$, an example is a group of commutative symmetric matrices, or equivalently, a group of the form $U\mathcal{D}U^{\mathsf{T}}$ where $U$ is a fixed orthogonal matrix and $\mathcal{D}$ is a diagonal group.
If $\mathcal{B}$ is not restricted to the form $\beta\beta_{0}\operatorname{I}_{\mathfrak{a}}$, there are more possibilities, we discuss an example in section 5. We will not be able to discuss an example where the bilinear form is not a trace form.
$\mathfrak{a}$ and $\mathfrak{a}_{{}_{\perp}}$ $\beta_{0}$ $\beta_{1}$ Examples
$(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\}$, $\mathfrak{a}_{\mathrm{sym}}=\{0\},\mathfrak{g}_{\mathrm{sym}}=\{0\}$ - ++ $\mathcal{A}\subset\mathcal{G}\subset\mathrm{SO}(n)$.
$(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\},\mathfrak{a}_{\mathrm{skew}}=\{0\}$ - - $\mathcal{D}=\mathcal{A}\subset\mathcal{G}=\mathrm{SO}(n)\times\mathcal{D}$.
$(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},\mathfrak{a}_{\mathrm{sym}}=\{0\}$ ++ ++ $\mathrm{SO}(n)=\mathcal{A}\subset\mathcal{G}=\operatorname{GL}(n)$.
$(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},\mathfrak{a}_{\mathrm{skew}}=\{0\},\mathfrak{g}_{\mathrm{skew}}=\{0\}$ ++ - $\mathcal{A}\subset\mathcal{G}=U\mathcal{D}U^{\mathsf{T}}$, $UU^{\mathsf{T}}=\operatorname{I}_{n}$.
Table 1. Summary of Riemannian left-invariant metrics obtained by deforming a transposable algebra $\mathfrak{g}$ using a transposable $\mathfrak{a}\subset\mathfrak{g}$. We assume $\mathcal{D}$ is a diagonal group.
5. The General Linear group with positive determinant
In this case, the group is $\mathcal{G}=\operatorname{GL}^{+}(n)$, the group of invertible matrices with positive determinant in $\mathbb{R}^{n\times n}$, with Lie algebra $\mathfrak{g}=\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$. We take $\mathfrak{a}=\mathfrak{so}(n)$, the Lie algebra of antisymmetric matrices. For $\mathtt{X}\in\mathfrak{g}$, we have $\mathcal{P}_{\mathfrak{a}}\mathtt{X}=\mathtt{X}_{\mathrm{skew}}$. We typically take $\beta_{0}=1$, and consider $\mathcal{B}=\beta\operatorname{I}_{\mathfrak{a}}$. For $\mathtt{X}\in\mathfrak{g}$, the metric becomes
(32) $\operatorname{Tr}\mathtt{X}^{2}-(\beta+1)\operatorname{Tr}\mathtt{X}_{\mathrm{skew}}\mathtt{X}_{\mathrm{skew}}=\operatorname{Tr}\mathtt{X}^{2}_{\mathrm{sym}}+\beta\operatorname{Tr}\mathtt{X}_{\mathrm{skew}}^{\mathsf{T}}\mathtt{X}_{\mathrm{skew}}=\lvert\mathtt{X}\rvert_{\mathcal{F}}^{2}+(\beta-1)\lvert\mathtt{X}_{\mathrm{skew}}\rvert_{\mathcal{F}}^{2}$
using the decomposition $\mathtt{X}=\mathtt{X}_{\mathrm{sym}}+\mathtt{X}_{\mathrm{skew}}$. We need $\beta>0$ for the metric to be Riemannian. When $\beta=1$, this is the Frobenius form, covered in [42]. Expanding the $\mathrm{skew}$ operator, theorem 2 implies
@@ 5. The General Linear group with positive determinant
The geodesic starting at $X$ with initial velocity $\xi=X\mathtt{V}$ is given by
(34) $\gamma(t)=X\exp(\frac{t}{2}\left((1-\beta)\mathtt{V}+(1+\beta)\mathtt{V}^{\mathsf{T}}\right))\exp(t(1+\beta)\mathtt{V}_{\mathrm{skew}}).$
@@
(35) $\mathtt{T}_{0}^{t}(\gamma)\eta=X\exp(\frac{t}{2}\left((1-\beta)\mathtt{V}+(1+\beta)\mathtt{V}^{\mathsf{T}}\right))\operatorname{expv}(t\mathrm{P},X^{-1}\eta)\exp(t(1+\beta)(\mathtt{V}_{\mathrm{skew}})),$
where the operator $\mathrm{P}=\mathrm{P}_{\mathtt{V}}$ is given by
(36) $\mathrm{P}_{\mathtt{V}}\mathtt{B}=\frac{1}{2}([\mathtt{B},\mathtt{V}]+(1+\beta)([\mathtt{V}_{\mathrm{skew}},\mathtt{B}]-[\mathtt{B}_{\mathrm{skew}},\mathtt{V}])).$
We illustrate theorem 1 by discussing the naturally reductive structure of the Frobenius metric in [42], corresponding to $\beta=1$. The metric eq. 18 for $\xi,\eta\in T_{X}\operatorname{GL}(n)$ is $\operatorname{Tr}(X^{-1}\xi(X^{-1}\eta)^{\mathsf{T}})$. The group $\tilde{\mathcal{G}}$ is $\operatorname{GL}(n)\times\mathrm{SO}(n)$, which acts on $\mathcal{G}=\operatorname{GL}(n)$ by $(X,A).Y=XYA^{\mathsf{T}}$, $X,Y\in\operatorname{GL}(n)$, $A\in\mathrm{SO}(n)$ (see also [24]). The lift of $\mathtt{X}\in\mathfrak{g}=\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$ in eq. 25 is $(\mathtt{X}^{\mathsf{T}},\mathtt{X}^{\mathsf{T}}-\mathtt{X})$, and the horizontal space $\tilde{\mathfrak{m}}$ consists of vectors of this form. The metric eq. 22 is
[…]
and evaluating the inner product on the lifts shows this is a submersion to $\operatorname{GL}(n)$ with the Frobenius metric $(\mathtt{X}_{1},\mathtt{X}_{2})=\operatorname{Tr}\mathtt{X}_{1}^{\mathsf{T}}\mathtt{X}_{2}$. The $\operatorname{ad}_{\tilde{\mathcal{H}}}-$invariance condition on $\tilde{\mathfrak{m}}$ implies $\operatorname{ad}_{\mathfrak{so}(n)}$-invariance on $\mathfrak{gl}(n)$, that means for $\mathtt{A}\in\mathfrak{so}(n),\mathtt{X}_{1},\mathtt{X}_{2}\in\mathfrak{gl}(n)$
$\operatorname{Tr}[\mathtt{A},\mathtt{X}_{1}]^{\mathsf{T}}\mathtt{X}_{2}+\operatorname{Tr}\mathtt{X}_{1}^{\mathsf{T}}[\mathtt{A},\mathtt{X}_{2}]=0,$
which also follows from $\operatorname{Tr}[\mathtt{A},\mathtt{X}_{1}]^{\mathsf{T}}\mathtt{X}_{2}=\operatorname{Tr}[\mathtt{X}_{1}^{\mathsf{T}},-\mathtt{A}]\mathtt{X}_{2}=-\operatorname{Tr}[\mathtt{A},\mathtt{X}_{2}]\mathtt{X}_{1}^{\mathsf{T}}$. For the naturally reductive condition eq. 9, the projection $\mathcal{P}_{\tilde{\mathfrak{m}}}$ is given by $\mathcal{P}_{\tilde{\mathfrak{m}}}(\mathtt{X},\mathtt{A})=(\mathtt{X}^{\mathsf{T}}+\mathtt{A},2\mathtt{A}+\mathtt{X}^{\mathsf{T}}-\mathtt{X})$ from eq. 20. Denote $\tilde{\mathtt{X}}\in\tilde{\mathfrak{m}}$ for the lift of $\mathtt{X}\in\mathfrak{gl}(n)$ then for $\mathtt{X}_{1},\mathtt{X}_{2}\in\mathfrak{gl}(n)$
$\mathcal{P}_{\tilde{\mathfrak{m}}}[\tilde{\mathtt{X}}_{1},\tilde{\mathtt{X}}_{2}]=\mathcal{P}_{\tilde{\mathfrak{m}}}([\mathtt{X}_{1}^{\mathsf{T}},\mathtt{X}_{2}^{\mathsf{T}}],[\mathtt{X}_{1}^{\mathsf{T}}-\mathtt{X}_{1},\mathtt{X}_{2}^{\mathsf{T}}-\mathtt{X}_{2}]),$
which is the lift of $\mathtt{X}_{1}^{\mathsf{T}},\mathtt{X}_{2}^{\mathsf{T}}]-[\mathtt{X}_{1}^{\mathsf{T}}-\mathtt{X}_{1},\mathtt{X}_{2}^{\mathsf{T}}-\mathtt{X}_{2}]=[\mathtt{X}_{1}^{\mathsf{T}},\mathtt{X}_{2}]+[\mathtt{X}_{1},\mathtt{X}_{2}^{\mathsf{T}}]-[\mathtt{X}_{1},\mathtt{X}_{2}]$. Thus, condition eq. 9, which follows from the quotient structure, would also follow from the following identity, which can be verified directly by manipulating trace identities
(38) $\operatorname{Tr}([\mathtt{X}_{1}^{\mathsf{T}},\mathtt{X}_{2}]+[\mathtt{X}_{1},\mathtt{X}_{2}^{\mathsf{T}}]-[\mathtt{X}_{1},\mathtt{X}_{2}])\mathtt{X}_{3}^{\mathsf{T}}=\operatorname{Tr}([\mathtt{X}_{3}^{\mathsf{T}},\mathtt{X}_{1}]+[\mathtt{X}_{3},\mathtt{X}_{1}^{\mathsf{T}}]-[\mathtt{X}_{3},\mathtt{X}_{1}])\mathtt{X}_{2}^{\mathsf{T}}.$
@@
Instead of taking $\mathfrak{a}=\mathfrak{so}(n)$, if we take $\mathfrak{a}=\mathfrak{so}(n)\oplus\mathbb{R}\operatorname{I}_{n}$, then the scalar operator $\beta\operatorname{I}_{\mathfrak{a}}$ violates the condition in item 3 of the list of possible choices of $\beta$ to produce a Riemannian metric. However, if we take instead the operator $\mathcal{B}$ which acts by multiplying $\mathfrak{so}(n)$ by $\beta$ and $\mathbb{R}\operatorname{I}_{n}$ by a negative constant $\mu$, that is for $\mathtt{A}\in\mathfrak{a}$
$\mathcal{B}\mathtt{A}=\beta(\mathtt{A}-\frac{1}{n}\operatorname{Tr}(\mathtt{A})\operatorname{I}_{n})-\frac{\mu}{n}\operatorname{Tr}(\mathtt{A})\operatorname{I}_{n}.$
Then the metric in section 4.3 is Riemannian, the formulas still hold, as $\mathbb{R}\operatorname{I}_{n}$ does not contribute to the Lie brackets. This is the metric in [24].
6. The Special Orthogonal group with a deformation metric
@@
In this case, $\mathcal{G}=\mathrm{SO}(n)$, the group of orthogonal matrices with determinant $1$. Its Lie algebra $\mathfrak{g}=\mathfrak{so}(n)$ is identified with the set of antisymmetric matrices. For $\mathfrak{a}$, we take the algebra of matrices where the top $d\times d$ block is antisymmetric and the remaining entries vanish, ($\mathfrak{a}=\begin{bmatrix}\mathfrak{so}(d)&0_{d\times(n-d)}\\ 0_{(n-d)\times d}&0_{(n-d)\times(n-d)}\end{bmatrix}$ graphically). Here, the projection $\mathcal{P}_{\mathfrak{a}}$ from $\mathbb{R}^{n\times n}$ maps a matrix $\mathtt{W}$ to the matrix $\mathcal{P}_{\mathfrak{a}}\mathtt{W}$ with the top $d\times d$ block is antisymmetrized from the corresponding block of $\mathtt{W}$, while the remaining entries vanish. The standard normalization takes $\beta_{0}=-\frac{1}{2}$. For $\mathcal{B}=\alpha\operatorname{I}_{\mathfrak{a}}$, we have
@@
At $X\in\mathrm{SO}(n)$, $\xi,\eta\in T_{X}\mathrm{SO}(n)$ where $\xi=X\begin{bmatrix}A_{\xi}&-B^{\mathsf{T}}_{\xi}\\ B_{\xi}&C_{\xi}\end{bmatrix}$, $\eta=X\begin{bmatrix}A_{\eta}&-B^{\mathsf{T}}_{\eta}\\ B_{\eta}&C_{\xi}\end{bmatrix}$, with $A_{\xi},A_{\eta}\in\mathfrak{so}(d),C_{\xi},C_{\eta}\in\mathfrak{so}(n-d)$, the metric eq. 18 with $\beta_{0}=-\frac{1}{2},\mathcal{B}=\alpha\operatorname{I}_{\mathfrak{a}}$ is given by
(39) $\langle\xi,\eta\rangle=\frac{1}{2}\operatorname{Tr}\xi^{\mathsf{T}}\eta+(\alpha-\frac{1}{2})\operatorname{Tr}(\mathcal{P}_{\mathfrak{a}}(\xi^{\mathsf{T}}X)\mathcal{P}_{\mathfrak{a}}(X^{\mathsf{T}}\eta))=\frac{1}{2}\operatorname{Tr}\xi^{\mathsf{T}}\eta+(\alpha-\frac{1}{2})\operatorname{Tr}A_{\xi}^{\mathsf{T}}A_{\eta}.$
@@
[…]
@@
[…]
@@
[…]
where the operator $\mathrm{P}=\mathrm{P}_{\mathtt{V}}$ for $\mathtt{V}=X^{\mathsf{T}}\xi$ is given in eq. 29.
This is the metric in [16] for the $\mathrm{SO}(n)$ identified with the Stiefel manifold $\mathrm{St}_{n,n}$.
7. A naturally reductive family of quotient metrics
We now introduce a deformation metric on a quotient manifold $\mathcal{G}/\mathcal{K}$ where $\mathcal{G}$ has a deformation metric by a subgroup $\mathcal{A}$. The resulting metric is also naturally reductive. The two standard metrics on the Stiefel manifold are instances of these metrics. Roughly, the requirement is the Lie algebra of $\mathcal{K}$ is both orthogonal and commuting with the Lie algebra $\mathfrak{a}$ of $\mathcal{A}$. The precise statement is in theorem 3 below.
Let us explain through an example of a space of block matrices. For $k<n$, let $\mathfrak{a}$ be a copy of $\mathbb{R}^{k\times k}$, considered as a subspace of $\mathcal{E}=\mathbb{R}^{n\times n}$ by identifying $\mathfrak{a}$ with the space of upper-left $k\times k$ block diagonal matrices, and $\mathfrak{a}_{{}_{\perp}}$ its Frobenius orthogonal complements. Hence, $\mathfrak{a}$ consists of matrices of the form $\begin{bmatrix}A&0\\ 0&0\end{bmatrix}$ with $A\in\mathbb{R}^{k\times k}$, while $\mathfrak{a}_{{}_{\perp}}$ consists of those of the form $\begin{bmatrix}0&B\\ C&D\end{bmatrix}$, with $D\in\mathbb{R}^{(n-k)\times(n-k)},B\in\mathbb{R}^{k\times(n-k)},C\in\mathbb{R}^{(n-k)\times k}$. Notice that $\mathfrak{a}$ and $\mathfrak{k}=\{\begin{bmatrix}0&0\\ 0&D\end{bmatrix}|D\in\mathbb{R}^{(n-k)\times(n-k)}\}$ are both commuting and orthogonal. This is the situation we are interested in.